------------------------------------------------------------ Date: Fri Dec 28 00:01:07 2001, Shift: 16:00-24:00 Supervisor: R. Debbe, with: Log book#11 pages 101 - 103 Shift Summary: ------------------------------------------------------------ 1. I arrived while the area was on restricted access. Up to now no beam has been injected yet. 2. No beams, but Bob and John fixed the 1458B mainframe, Flemming had fb3 back to life sometime ago. BFS is thus ready to be added to the data stream. All magnets are set at B polarity. 3. All PMTs are basking happily in their dark current, TPCs and drift ch. are in park mode. No problems, no achievements (do we have a pulse?) ------------------------------------------------------------
This archive was generated by hypermail 2b30 : Fri Dec 28 2001 - 00:01:38 EST